# Sparse Representations Improve Adversarial Robustness of Neural Network Classifiers

### v1

@@ 1 Introduction
In this paper we explore dimensionality reduction as a defense and ask whether enforcing sparsity in the projection improves robustness. Specifically, we investigate Sparse Principal Component Analysis (SPCA, spca ) as a data-adapted linear front-end in place of standard PCA pca . Our hypothesis builds on the view that adversarial examples exploit high-dimensional, non-robust features that correlate with labels in training data but fail to generalize under small perturbations. By restricting each component to depend on only a small subset of input dimensions, SPCA may filter out these non-robust directions and emphasise more stable structure.
Beyond empirical evaluation, we provide a theoretical account of why sparsity helps. We derive exact robustness certificates for linear heads operating on SPCA features. In the $\ell_{\infty}$ and $\ell_{2}$ settings—both binary and multiclass—the prediction is provably invariant whenever the clean margin exceeds a term proportional to $\|W^{\top}u\|_{1}$ or $\|W^{\top}u\|_{2}$, respectively, where $W$ is the projection and $u$ are head weights. Because SPCA drives many entries of $W$ to zero and shrinks its column norms, these dual norms decrease, enlarging the certified radius. For general non-linear heads, we bound the end-to-end sensitivity via a Lipschitz composition $\|C_{\phi}\circ W\|_{p\to 2}\leq L_{C}\|W\|_{p\to 2}$, showing that sparsity in $W$ tightens operator-norm bounds and thus reduces the model’s worst-case input sensitivity.
@@ 2 Related Work
A large body of work investigates why modern classifiers are brittle and how to defend them. One perspective attributes adversarial vulnerability to the exploitation of non-robust features, i.e., high-dimensional but brittle correlations that drive accuracy on clean data while being easy to manipulate ilyas2019adversarialnotbugs . This motivates defenses that either (i) project inputs onto lower-complexity representations to discard non-robust directions, or (ii) change the classifier family altogether (e.g., generative classifiers) so that decisions align with data likelihood. Our work follows the first route via linear dimensionality reduction with sparsity, and we position it relative to both lines below.
@@
Prior defenses reduce the adversary’s search space by transforming inputs before classification. Representative examples include bit-depth reduction, JPEG compression, total-variation minimisation, and image quilting guo2018countering , as well as feature squeezing (bit-depth and spatial smoothing) used for detection xu2017featureSqueezing . These methods share the spirit of dimensionality/complexity reduction but typically operate as non-learned image transforms. In contrast, we learn a data-adapted linear projection and ask whether imposing sparsity on this projection improves robustness over standard PCA.
@@
High-dimensional models like deep neural networks are famously vulnerable to small adversarial perturbations. A common defense strategy is to reduce input dimensionality or project inputs onto a lower-dimensional “manifold” of normal data, with the goal of discarding adversarial noise. Prior work has shown promise for such approaches. For example, bhagoji2017enhancingrobustnessmachinelearning explored data transformations including PCA as a preprocessing defense, reporting that projecting inputs onto a subspace of top principal components can significantly increase the effort required for a successful attack. This supports the intuition that many adversarial perturbations lie in directions of input space that can be truncated or removed through dimensionality reduction.
@@
Researchers have also analysed adversarial examples through the lens of principal components. Hendrycks and Gimpel hendrycks2017earlymethodsdetectingadversarial observed that adversarial images exhibit different PCA spectra than genuine images. In particular, adversarial inputs tend to have abnormally large variance along some principal directions. jere2019principalcomponentpropertiesadversarial extended this analysis across multiple architectures, showing that small adversarial perturbations consistently alter an input’s projection on principal components in characteristic ways. Such findings motivated the use of PCA as a defensive filter; for instance, bhagoji2017enhancingrobustnessmachinelearning demonstrated that removing low-variance PCA components of input data can improve robustness against $\ell_{2}$ and $\ell_{\infty}$ norm attacks, forcing attackers to exert significantly more distortion to fool the classifier.
@@
Sparse Principal Component Analysis is a variant of PCA that seeks a projection onto principal components with sparse support (i.e. each component is a linear combination of only a small subset of original features) spca . By selecting a subset of important features, SPCA can yield more interpretable and possibly more robust representations, since spurious noise spread across many features might be excluded. Despite SPCA’s popularity in high-dimensional data analysis, its use as a defense against adversarial attacks remains largely unexplored. To our knowledge, no prior empirical study has focused on applying SPCA to improve classifier resistance to adversarial perturbations in image recognition tasks. Related work has predominantly centred on standard PCA or non-sparse transformations, as discussed above.
@@
There are a few recent developments that hint at the potential of SPCA-like approaches for robustness. awasthi2020estimatingprincipalcomponentsadversarial theoretically formulated a “robust PCA” objective for finding low-dimensional data representations that are maximally stable to bounded adversarial perturbations. Interestingly, their formulation includes a sparse PCA criterion as a special case, and they devise algorithms to compute a robust subspace with provable guarantees. A complementary theoretical analysis of PCA’s susceptibility to worst-case perturbations was given by li2020intheadversarialrobustnesspca , who studied optimal modification strategies that maximize the distance between subspaces learned from clean versus perturbed data. Their results provide formal insights into how adversaries can exploit the structure of the data matrix to destabilise PCA projections, especially when the adversary can manipulate the entire dataset. dorsi2020sparsepcaalgorithmsadversarial likewise study the problem of identifying the top sparse principal components in a dataset that has been adversarially corrupted, providing the first algorithms with certified resilience to small adversarial changes in the input covariance. These works are theoretical and focus on algorithmic recovery guarantees rather than end-to-end classifier defense, but they underscore the intuition that enforcing sparsity in feature projections can enhance robustness.
@@
Orthogonal to dimensionality-reduction defenses, li2019aregenerative propose deep generative classifiers that model $p(x,y)$ and use likelihood for both prediction and attack detection. Their experiments cover MNIST and a CIFAR-binary task constructed from CIFAR-10 (airplane vs. frog), and report improved robustness against common attacks. Conceptually, their robustness arises from rejecting off-manifold inputs via low likelihood, whereas our approach aims to prevent vulnerability by discarding non-robust directions through a sparse linear projection learned from data. The two lines are complementary: SPCA can be used as a front-end to generative or discriminative models, and our evaluation protocol (MNIST/CIFAR-binary; FGSM/PGD/MIM; $\ell_{\infty}$ and $\ell_{2}$) mirrors li2019aregenerative in spirit but targets a different mechanism.
The view that adversaries exploit non-robust features ilyas2019adversarialnotbugs provides a rationale for favoring sparse representations that emphasise a small, stable subset of input dimensions. Robust-subspace formulations for PCA provide algorithmic backing for this intuition awasthi2020estimatingprincipalcomponentsadversarial , while certified approaches like randomized smoothing cohen2019certifiedsmoothing pursue orthogonal guarantees. Our findings suggest that enforcing sparsity already yields meaningful empirical robustness improvements and can potentially complement certified or generative defenses.
@@ 3 Problem Statement
Our aim is to compare the robustness of two classifiers: one built upon the PCA‑based feature extractor $W_{\mathrm{PCA}}$, and another on the SPCA‑based extractor $W_{\mathrm{SPCA}}$. Specifically, we wish to investigate whether enforcing sparsity in the linear transformation $\Phi$ leads to improved adversarial robustness. Hence, we study $\mathrm{Robustness}(h_{\mathrm{PCA}},\varepsilon)\quad\text{vs.}\quad\mathrm{Robustness}(h_{\mathrm{SPCA}},\varepsilon)$, for various perturbation magnitudes $\varepsilon$, where
@@ 4.2 Certified Robustness for Binary Linear Heads
Theorem 1 (Exact $\ell_{\infty}$ robustness certificate).
@@
Proof.
For any such $\delta$, $\big|u^{\top}W\delta\big|\leq\varepsilon\,\|W^{\top}u\|_{1}$ by $\ell_{\infty}$-$\ell_{1}$ duality. Hence the logit $y(u^{\top}W(x+\delta)+b)$ remains positive if $m(x)>\varepsilon\|W^{\top}u\|_{1}$. ∎
Theorem 2 (Exact $\ell_{2}$ robustness certificate).
@@
Proof.
By Cauchy-Schwarz, $\big|u^{\top}W\delta\big|\leq\|W^{\top}u\|_{2}\,\|\delta\|_{2}\leq\varepsilon\,\|W^{\top}u\|_{2}$. ∎
@@ 4.3 Multiclass Linear Heads
Theorem 3 (Multiclass $\ell_{\infty}$ certificate).
@@
Proof.
For each competitor $k$, the perturbed logit difference equals $(u_{k^{\star}}-u_{k})^{\top}W(x+\delta)+(b_{k^{\star}}-b_{k})$, whose change due to $\delta$ is bounded in magnitude by $\varepsilon\|W^{\top}(u_{k^{\star}}-u_{k})\|_{1}$ by duality. If the clean margin exceeds this bound for all $k$, all pairwise differences remain positive. ∎
Theorem 4 (Multiclass $\ell_{2}$ certificate).
@@
Proof.
As in Theorem 2, by Cauchy-Schwarz the change in each pairwise logit difference is at most $\varepsilon\,\|W^{\top}(u_{k^{\star}}-u_{k})\|_{2}$. ∎
@@ 4.5 How Sparsity in $W$ Improves Certificates
Lemma 1 (Dual-norm control via column norms).
For any $u\in\mathbb{R}^{r}$, $\|W^{\top}u\|_{1}\leq\|u\|_{2}\sum_{j=1}^{D}\|w_{j}\|_{2}$ and $\|W^{\top}u\|_{2}\leq\|u\|_{2}\,\|W\|_{2\to 2}.$
Proof.
Write $(W^{\top}u)_{j}=u^{\top}w_{j}$. Then $\|W^{\top}u\|_{1}=\sum_{j=1}^{D}|u^{\top}w_{j}|\leq\sum_{j=1}^{D}\|u\|_{2}\,\|w_{j}\|_{2}=\|u\|_{2}\sum_{j=1}^{D}\|w_{j}\|_{2}$ by Cauchy-Schwarz. The $\ell_{2}$ bound is the definition of the operator norm $\|W\|_{2\to 2}$ applied to $u$. ∎
@@ 5.1 Datasets and Preprocessing
The first dataset is the widely used MNIST handwritten digits benchmark. It contains 60,000 training images and 10,000 test images of grayscale digits normalized to $28\times 28$ pixels deng2012mnist . We follow the conventional split, using the entire training set to learn the PCA or SPCA projection matrix and to train a downstream classifier, and evaluating robustness on the 10,000‑image test set. After flattening each image into a 784‑dimensional vector and mean‑centring, we project onto $r$ principal components. For SPCA we use a sparsity parameter that retains approximately $5\%$ of the weights in each component.
@@
Our second dataset is a binary version of CIFAR‑10 krizhevsky2009learningcifar inspired by li2019aregenerative . CIFAR‑10 contains 60,000 $32\times 32$ color images in 10 classes, with 50,000 training examples and 10,000 test examples. To create the binary variant we extract only the classes “airplane” and “frog” from the original dataset, yielding 6,000 airplane images and 6,000 frog images. We split these into 10,000 training examples and 2,000 test examples. Each image is converted to grayscale, flattened into a 1024‑dimensional vector, mean‑centred, and projected onto $r$ principal components with PCA or SPCA.
@@ 5.3 Attack Models
FGSM is a one‑step, white‑box attack that perturbs an input in the direction of the sign of the loss gradient fgsm . An adversarial example $x^{\prime}$ is generated by $x^{\prime}=x+\varepsilon\cdot\operatorname{sign}(\nabla_{x}\ell(h(x),y))$, where $\varepsilon$ is the attack strength. Despite its simplicity, FGSM effectively reduces classification accuracy even for very small $\varepsilon$.
@@
PGD is a multi‑step extension of FGSM that iteratively takes gradient steps and projects back onto the allowed perturbation set pgd . For perturbation budget $\varepsilon$ and step size $\alpha$ we perform $T$ iterations:
@@
MIM augments iterative gradient attacks with a momentum term to stabilise updates and improve transferability mim . At each iteration the gradient is accumulated into a velocity vector $g^{(t)}$ and normalized. The update rule is $g^{(t+1)}=\mu\cdot g^{(t)}+\frac{\nabla_{x^{(t)}}\ell(h(x^{(t)}),y)}{\|\nabla_{x^{(t)}}\ell(h(x^{(t)}),y)\|_{1}}$, and the adversarial example is updated as $x^{(t+1)}=x^{(t)}+\alpha\cdot\operatorname{sign}(g^{(t+1)})$ followed by projection. We use a momentum parameter $\mu=1.0$ and step size $\alpha=\varepsilon/5$ for 20 iterations.
@@
The Square Attack is a score‑based black‑box attack that queries only the model’s output probabilities and randomly perturbs square regions of the image until misclassification occurs andriushchenko2020squareattackqueryefficientblackbox . The method is query‑efficient and circumvents gradient masking by not requiring access to model gradients. We adopt the implementation of andriushchenko2020squareattackqueryefficientblackbox with 5,000 queries and report the success rate as a function of $\varepsilon$.

### v2

@@ Abstract
Keywords: adversarial robustness, sparse representations, neural network classifiers, robust feature learning, regularization
@@ 1 Introduction
In this paper we explore dimensionality reduction as a defense and ask whether enforcing sparsity in the projection improves robustness. Specifically, we investigate Sparse Principal Component Analysis (SPCA) from Zou et al. (2006) as a data-adapted linear front-end in place of standard PCA (Wold et al., 1987). Our hypothesis builds on the view that adversarial examples exploit high-dimensional, non-robust features that correlate with labels in training data but fail to generalize under small perturbations. By restricting each component to depend on only a small subset of input dimensions, SPCA may filter out these non-robust directions and emphasise more stable structure.
Beyond empirical evaluation, we provide a theoretical account of why sparsity helps. We derive exact robustness certificates for linear heads operating on SPCA features. In the $\ell_{\infty}$ and $\ell_{2}$ settings, both binary and multiclass, the prediction is provably invariant whenever the clean margin exceeds a term proportional to $\|W^{\top}u\|_{1}$ or $\|W^{\top}u\|_{2}$, respectively, where $W$ is the projection and $u$ are head weights. Because SPCA drives many entries of $W$ to zero and shrinks its column norms, these dual norms decrease, enlarging the certified radius. For general non-linear heads, we bound the end-to-end sensitivity via a Lipschitz composition $\|C_{\phi}\circ W\|_{p\to 2}\leq L_{C}\|W\|_{p\to 2}$, showing that sparsity in $W$ tightens operator-norm bounds and thus reduces the model’s worst-case input sensitivity.
@@ 2 Related Work
A large body of work investigates why modern classifiers are brittle and how to defend them. One perspective attributes adversarial vulnerability to the exploitation of non-robust features, i.e., high-dimensional but brittle correlations that drive accuracy on clean data while being easy to manipulate (Ilyas et al., 2019). This motivates defenses that either (i) project inputs onto lower-complexity representations to discard non-robust directions, or (ii) change the classifier family altogether (e.g., generative classifiers) so that decisions align with data likelihood. Our work follows the first route via linear dimensionality reduction with sparsity, and we position it relative to both lines below.
@@
Prior defenses reduce the adversary’s search space by transforming inputs before classification. Representative examples include bit-depth reduction, JPEG compression, total-variation minimization, and image quilting (Guo et al., 2018), as well as feature squeezing (bit-depth and spatial smoothing) used for detection (Xu et al., 2017). These methods share the spirit of dimensionality/complexity reduction but typically operate as non-learned image transforms. In contrast, we learn a data-adapted linear projection and ask whether imposing sparsity on this projection improves robustness over standard PCA.
@@
High-dimensional models like deep neural networks are famously vulnerable to small adversarial perturbations. A common defense strategy is to reduce input dimensionality or project inputs onto a lower-dimensional “manifold” of normal data, with the goal of discarding adversarial noise. Prior work has shown promise for such approaches. For example, Bhagoji et al. (2017) explored data transformations including PCA as a preprocessing defense, reporting that projecting inputs onto a subspace of top principal components can significantly increase the effort required for a successful attack. This supports the intuition that many adversarial perturbations lie in directions of input space that can be truncated or removed through dimensionality reduction.
@@
Researchers have also analysed adversarial examples through the lens of principal components. Hendrycks and Gimpel (2017) observed that adversarial images exhibit different PCA spectra than genuine images. In particular, adversarial inputs tend to have abnormally large variance along some principal directions. Jere et al. (2019) extended this analysis across multiple architectures, showing that small adversarial perturbations consistently alter an input’s projection on principal components in characteristic ways. Such findings motivated the use of PCA as a defensive filter; for instance, Bhagoji et al. (2017) demonstrated that removing low-variance PCA components of input data can improve robustness against $\ell_{2}$ and $\ell_{\infty}$ norm attacks, forcing attackers to exert significantly more distortion to fool the classifier.
@@
Sparse Principal Component Analysis is a variant of PCA that seeks a projection onto principal components with sparse support, i.e. each component is a linear combination of only a small subset of original features (Zou et al., 2006). By selecting a subset of important features, SPCA can yield more interpretable and possibly more robust representations, since spurious noise spread across many features might be excluded. Despite SPCA’s popularity in high-dimensional data analysis, its use as a defense against adversarial attacks remains largely unexplored. To our knowledge, no prior empirical study has focused on applying SPCA to improve classifier resistance to adversarial perturbations in image recognition tasks. Related work has predominantly centred on standard PCA or non-sparse transformations, as discussed above.
@@
There are a few recent developments that hint at the potential of SPCA-like approaches for robustness. Awasthi et al. (2020) theoretically formulated a “robust PCA” objective for finding low-dimensional data representations that are maximally stable to bounded adversarial perturbations. Interestingly, their formulation includes a sparse PCA criterion as a special case, and they devise algorithms to compute a robust subspace with provable guarantees. A complementary theoretical analysis of PCA’s susceptibility to worst-case perturbations was given by Li et al. (2020), who studied optimal modification strategies that maximize the distance between subspaces learned from clean versus perturbed data. Their results provide formal insights into how adversaries can exploit the structure of the data matrix to destabilise PCA projections, especially when the adversary can manipulate the entire dataset. d’Orsi et al. (2020) likewise study the problem of identifying the top sparse principal components in a dataset that has been adversarially corrupted, providing the first algorithms with certified resilience to small adversarial changes in the input covariance. These works are theoretical and focus on algorithmic recovery guarantees rather than end-to-end classifier defense, but they underscore the intuition that enforcing sparsity in feature projections can enhance robustness.
@@
Orthogonal to dimensionality-reduction defenses, Li et al. (2019) propose deep generative classifiers that model $p(x,y)$ and use likelihood for both prediction and attack detection. Their experiments cover MNIST and a CIFAR-binary task constructed from CIFAR-10 (airplane vs. frog), and report improved robustness against common attacks. Conceptually, their robustness arises from rejecting off-manifold inputs via low likelihood, whereas our approach aims to prevent vulnerability by discarding non-robust directions through a sparse linear projection learned from data. The two lines are complementary: SPCA can be used as a front-end to generative or discriminative models, and our evaluation protocol (MNIST/CIFAR-binary; FGSM/PGD/MIM; $\ell_{\infty}$ and $\ell_{2}$) mirrors Li et al. (2019) in spirit but targets a different mechanism.
The view that adversaries exploit non-robust features (Ilyas et al., 2019) provides a rationale for favoring sparse representations that emphasise a small, stable subset of input dimensions. Robust-subspace formulations for PCA provide algorithmic backing for this intuition (Awasthi et al., 2020), while certified approaches like randomized smoothing (Cohen et al., 2019) pursue orthogonal guarantees. Our findings suggest that enforcing sparsity already yields meaningful empirical robustness improvements and can potentially complement certified or generative defenses.
@@ 3 Problem Statement
Our aim is to compare the robustness of two classifiers: one built upon the PCA‑based feature extractor $W_{\mathrm{PCA}}$, and another on the SPCA‑based extractor $W_{\mathrm{SPCA}}$. Specifically, we wish to investigate whether enforcing sparsity in the linear transformation $\Phi$ leads to improved adversarial robustness. Hence, we study $\mathrm{Robustness}(h_{\mathrm{PCA}},\varepsilon)$ vs. $\mathrm{Robustness}(h_{\mathrm{SPCA}},\varepsilon)$, for various perturbation magnitudes $\varepsilon$, where
@@ 4.2 Certified Robustness for Binary Linear Heads
Theorem 1 (Exact $\ell_{\infty}$ robustness certificate)
@@
Proof For any such $\delta$, $\big|u^{\top}W\delta\big|\leq\varepsilon\,\|W^{\top}u\|_{1}$ by $\ell_{\infty}$-$\ell_{1}$ duality. Hence the logit $y(u^{\top}W(x+\delta)+b)$ remains positive if $m(x)>\varepsilon\|W^{\top}u\|_{1}$.
Theorem 2 (Exact $\ell_{2}$ robustness certificate)
@@
Proof By Cauchy-Schwarz, $\big|u^{\top}W\delta\big|\leq\|W^{\top}u\|_{2}\,\|\delta\|_{2}\leq\varepsilon\,\|W^{\top}u\|_{2}$.
@@ 4.3 Multiclass Linear Heads
Theorem 3 (Multiclass $\ell_{\infty}$ certificate)
@@
Proof For each competitor $k$, the perturbed logit difference equals $(u_{k^{\star}}-u_{k})^{\top}W(x+\delta)+(b_{k^{\star}}-b_{k})$, whose change due to $\delta$ is bounded in magnitude by $\varepsilon\|W^{\top}(u_{k^{\star}}-u_{k})\|_{1}$ by duality. If the clean margin exceeds this bound for all $k$, all pairwise differences remain positive.
Theorem 4 (Multiclass $\ell_{2}$ certificate)
@@
Proof As in Theorem 2, by Cauchy-Schwarz the change in each pairwise logit difference is at most $\varepsilon\,\|W^{\top}(u_{k^{\star}}-u_{k})\|_{2}$.
@@ 4.5 How Sparsity in $W$ Improves Certificates
Lemma 5 (Dual-norm control via column norms)
For any $u\in\mathbb{R}^{r}$,
$\|W^{\top}u\|_{1}\leq\|u\|_{2}\sum_{j=1}^{D}\|w_{j}\|_{2}$ and $\|W^{\top}u\|_{2}\leq\|u\|_{2}\,\|W\|_{2\to 2}.$
Proof Write $(W^{\top}u)_{j}=u^{\top}w_{j}$. Then $\|W^{\top}u\|_{1}=\sum_{j=1}^{D}|u^{\top}w_{j}|\leq\sum_{j=1}^{D}\|u\|_{2}\,\|w_{j}\|_{2}=\|u\|_{2}\sum_{j=1}^{D}\|w_{j}\|_{2}$ by Cauchy-Schwarz. The $\ell_{2}$ bound is the definition of the operator norm $\|W\|_{2\to 2}$ applied to $u$.
@@ 5.1 Datasets and Preprocessing
The first dataset is the widely used MNIST handwritten digits benchmark. It contains 60,000 training images and 10,000 test images of grayscale digits normalized to $28\times 28$ pixels (Deng, 2012). We follow the conventional split, using the entire training set to learn the PCA or SPCA projection matrix and to train a downstream classifier, and evaluating robustness on the 10,000‑image test set. After flattening each image into a 784‑dimensional vector and mean‑centring, we project onto $r$ principal components. For SPCA we use a sparsity parameter that retains approximately $5\%$ of the weights in each component.
@@
Our second dataset is a binary version of CIFAR‑10 (Krizhevsky, 2009) inspired by Li et al. (2019). CIFAR‑10 contains 60,000 $32\times 32$ color images in 10 classes, with 50,000 training examples and 10,000 test examples. To create the binary variant we extract only the classes “airplane” and “frog” from the original dataset, yielding 6,000 airplane images and 6,000 frog images. We split these into 10,000 training examples and 2,000 test examples. Each image is converted to grayscale, flattened into a 1024‑dimensional vector, mean‑centred, and projected onto $r$ principal components with PCA or SPCA.
@@ 5.3 Attack Models
FGSM is a one‑step, white‑box attack that perturbs an input in the direction of the sign of the loss gradient (Goodfellow et al., 2015). An adversarial example $x^{\prime}$ is generated by $x^{\prime}=x+\varepsilon\cdot\operatorname{sign}(\nabla_{x}\ell(h(x),y))$, where $\varepsilon$ is the attack strength. Despite its simplicity, FGSM effectively reduces classification accuracy even for very small $\varepsilon$.
@@
PGD is a multi‑step extension of FGSM that iteratively takes gradient steps and projects back onto the allowed perturbation set (Madry et al., 2019). For perturbation budget $\varepsilon$ and step size $\alpha$ we perform $T$ iterations:
@@
MIM augments iterative gradient attacks with a momentum term to stabilise updates and improve transferability (Dong et al., 2018). At each iteration the gradient is accumulated into a velocity vector $g^{(t)}$ and normalized. The update rule is $g^{(t+1)}=\mu\cdot g^{(t)}+\frac{\nabla_{x^{(t)}}\ell(h(x^{(t)}),y)}{\|\nabla_{x^{(t)}}\ell(h(x^{(t)}),y)\|_{1}}$, and the adversarial example is updated as $x^{(t+1)}=x^{(t)}+\alpha\cdot\operatorname{sign}(g^{(t+1)})$ followed by projection. We use a momentum parameter $\mu=1.0$ and step size $\alpha=\varepsilon/5$ for 20 iterations.
@@
The Square Attack is a score‑based black‑box attack that queries only the model’s output probabilities and randomly perturbs square regions of the image until misclassification occurs (Andriushchenko et al., 2020). The method is query‑efficient and circumvents gradient masking by not requiring access to model gradients. We adopt the implementation of Andriushchenko et al. (2020) with 5,000 queries and report the success rate as a function of $\varepsilon$.